Simplify and express each of the following in exponential form
25^4 ÷ 5^3

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(25)^4can be written as = (5 × 5)^4
= (5^2)^4 ... [∵(a^m)^n = a^mn ]

(5^2)^4 can be written as = (5)^2 x 4
= 5^8
Then
5^8÷5^3
5^8-3 ... [∵a^m÷a^n = a^m-n]
5^5

Thanks